ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 187, 2012

Abardia, J., E. Gallego and G. Solanes The Gauss–Bonnet theorem and Crofton-type formulas in complex space forms

287

Albers, P. and U. Frauenfelder On a theorem by Ekeland–Hofer  485

Andreatta, F. Coleman–Oort’s conjecture for degenerate irreducible curves

231

Bartholdi, L., O. Siegenthaler and P. Zalesskii The congruence subgroup problem for branch groups

419

Blair, D. E. and A. Mihai Homogeneity and local symmetry of complex (κ, μ)-spaces

451

Dochtermann, A. and C. Schultz Topology of Hom complexes and test graphs for bounding chromatic number

371

Douglas, R. G., G. Misra and J. Sarkar Contractive Hilbert modules and their dilations

141

Flores, R. J. On the idempotency of some composite functors  81

Frauenfelder, U. See P. Albers Gallego, E. See J. Abardia Garban, C., S. Rohde and O. Schramm Continuity of the SLE

trace in simply connected domains

23

Gogolev, A. How typical are pathological foliations in partially hyperbolic dynamics: an example

493

Helm, D. A geometric Jacquet–Langlands correspondence for U(2) Shimura varieties

37

Keller, C. and M. A. Perles On the smallest sets blocking simple perfect matchings in a convex geometric graph

465

Kochloukova, D. H., C. Martínez-Pérez and B. E. A. Nucinkis Fixed points of finite groups acting on generalised Thompson groups

167

Král’, D., O. Serra and L. Vena A removal lemma for systems of linear equations over finite fields

193

Laghribi, A. Autour des formes quadratiques indécomposables de l’image de l’homomorphisme W (F) ⟶ W (F (ϕ))

1

Martínez-Pérez, C. See D. H. Kochloukova Mihai, A. See D. E. Blair Misra, G. See R. G. Douglas Murase, A. and H. Narita Fourier expansion of Arakawa lifting I:

An explicit formula and examples of non-vanishing lifts 317

Muskat, J. and R. Pries Alternating group covers of the affine line  117

Narita, H. See A. Murase Nazarov, F., Y. Peres and P. Shmerkin Convolutions of Cantor

measures without resonance

93

Newelski, L. Bounded orbits and measures on a group  209

Nucinkis, B. E. A. See D. H. Kochloukova Peres, Y. See F. Nazarov Perles, M. A. See C. Keller Pries, R. See J. Muskat Rohde, S. See C. Garban Sarkar, J. See R. G. Douglas Schramm, O. See C. Garban Schultz, C. See A. Dochtermann Serra, O. See D. Král’ Shmerkin, P. See F. Nazarov Siegenthaler, O. See L. Bartholdi Solanes, G. See J. Abardia Vena, L. See D. Král’ Zalesskii, P. See L. Bartholdi